OpenStudy (anonymous):

Factor the polynomials : 12x^2-26x-10

myininaya (myininaya):

\[12x^2-26x-10 \] \[2 \cdot 6x^2-2 \cdot 13x-2 \cdot 5\] \[2(6x^2-13x-5)\]

myininaya (myininaya):

now see if there are two factors of 6(-5) that have product 6(-5) and have sum -13 6(-5)=-30=2(-15) -13=2-15

myininaya (myininaya):

\[6x^2-15x+2x-5\] \[3x(2x-5)+1(2x-5)\] \[(2x-5)(3x+1)\] ---------------- \[2(6x^2-13x-5)=2(2x-5)(3x+1)\]